Grade 11Physical Chemistry

If an electron is continuously accelerated in vacuum through a potential difference, it`s wavelength decreases by 1%.The change in kinetic energy is nearly?

Rahul
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
 \left (\frac{dK.E.}{K.E.} \right )\times 100 =2\left (\frac{d\lambda.}{\lambda} \right )\times 100=20%So,\left (\frac{d\lambda}{\lambda} \right ) =\frac{-1}{2}}\left (\frac{dK.E.}{K.E} \right )So, we get:\lambda =\frac{h}{\sqrt{2mK.E}}=\frac{h}{\sqrt{2m}}\left ( K.E \right )^{-\frac{1}{2}}Dear student,We have the equation being given by:
Regards
Arun (askIITians forum expert)